How does a buffer resist change in pH upon additio
Answer #1
Concepts and reason

Buffer solution is a solution which maintains its pH value on addition of small amount of strong acid or strong base into it. It is a solution which resists the change in pH.

Fundamentals

Acidic buffer solutions are the solutions which have a pH less than 7. These solutions are made from weak acid and its salt with a strong base.

Basic buffer solutions are the solutions which have a pH greater than 7. These solutions are made from weak base and its salt with a strong acid.

There is no significant change in pH of the buffer solution by the addition of small amount of strong acid or base to it.

Consider a buffer solution that consist of a weak acid HA and its conjugate base A{{\rm{A}}^ - } . When a strong acid is added to buffer solution then the conjugate base, A{{\rm{A}}^ - } present in buffer solution takes its hydronium ion converting it into weak acid of conjugate base and water.

H3O+(aq)+A(aq)H2O(l)+HA(aq){{\rm{H}}_3}{{\rm{O}}^ + }\left( {aq} \right) + {{\rm{A}}^ - }\left( {aq} \right) \to {{\rm{H}}_2}{\rm{O}}\left( l \right) + {\rm{HA}}\left( {aq} \right)

Due to this, amount of weak acid increases and amount of conjugate base decreases.

By adding strong acid in buffer solution, amount of weak acid increases. So, there will be a little change in pH.

Therefore, the correct option is as follows:

The strong acid reacts with the weak base in the buffer to form a weak acid, which produces few H+{{\rm{H}}^ + } ions in solution and therefore only a little change in pH.

Ans:

The strong acid reacts with the weak base in the buffer to form a weak acid, which produces few H+{{\rm{H}}^ + } ions in solution and therefore only a little change in pH.